OpenStudy (anonymous):

easy multiple choice question! trig find the exact value of sin(x)+cos(x)-tan(x0/sec(x)+csc(x)-cot(x) when tan(x)=-4/3 is the answer a) -23/5 and -34/35 b) -23/5 and 34/35 c) 23/5 and 34/35 d) 23/5 and -34/35

OpenStudy (anonymous):

tan (y/x) ? so y=-4 and x=3 and just plug it in?

OpenStudy (anonymous):

x^2+y^2=r^2 to find r

OpenStudy (anonymous):

sin (y/r) +cos(x/r)-tan..... and plug in -4 for y and 3 for x

OpenStudy (anonymous):

is it a, b ,c or d???

OpenStudy (anonymous):

what's tan(x0/..... ) is that a zero or theta?

OpenStudy (anonymous):

sorry it should be tan(x) no zero